Mason Manufacturing is contemplating offering a new $60 million bond issue to replace an outstanding $60 million bond issue. The firm wishes to do this to take advantage of the decline in interest rates that has occurred since the initial bond issuance. The old and new bonds are described below. The firm is in the 30 per cent tax bracket.

Old bonds. The outstanding bonds have a $1,000 par value and a 7.5 per cent coupon interest rate.

They were issued three years ago with a 15-year maturity. They were initially sold for their par value of $1,000, and the firm incurred $345,000 in floatation costs. They are callable at $1,070.

New bonds. The new bonds would have a $1,000 par value, a 6 per cent coupon interest rate, and a 12-year maturity. They could be sold at their par value. The floatation cost of the new bonds would be $360,000. The firm expects to have two months of overlapping interest.

Required:

a. Calculate the initial investment that is required to call the old bonds & issue the new bonds.

b. Calculate the annual cashflow savings, if any, expected form the proposed bond refunding decision.

c. Would you recommend the proposed refunding? Why? Show your calculations clearly to explain your answer.
Answer #1

Given infromation,

Old bonds issued at $1000 with 7.5% coupon rate and 15 years to maturity = $60,000,000

Flotation cost of old bonds = $345,000

Callable value of bonds = $1070

New bonds issued at $1000 with 6% coupon rate and 12 years to maturity = $60,000,000

flotation cost of new bonds = $360,000

Firm expects to have over lapping interest for two months.

Computation of intial investment required to call old bonds and issue new bonds

Number of bonds = $60,000,000 \div $1000 = 60,000 bonds

Amount to be paid on old bonds = $1070 \times 60,000 = $64,200,000

Amount to be recieved on issue of new bonds = $1000 \times 60,000 = $60,000,000.

a) Call Premium

Premium on redemption of old bonds = $4,200,000

Less: Tax on premium @30% = ($1,260,000)

After tax cost of premium = $2,940,000

b) Flotation cost

Flotation cost on issue of new bonds = $360,000

c) Overlapping interest

Overlapping interest for two months = $60 million \times 7.5% \times 2 \div 12 = $750,000

Less: Tax @30% = 750,000 \times 30/100 = ($225000)

  After tax overlapping Interest cost = $525,000

d) Tax saving on unamortized flotation cost

Cost of old bond [$345,000 \times 12/15 \times 0.30] = $82,800

Total amount of initial investment = [a+b+c+d] = $3,907,800

Computation of annual cash savings

a) Old bond

Interest on old bonds = $60,000,000 \times 7.5% = $4,500,000

Less: Tax @ 30% = ($1,350,000)

After tax interest on old bonds = $3,150,000

Tax saving on amortisation of flotation cost = $345,000 \div 15 \times 0.3 = ($6900)

Annual after tax cost payment under old bond (a) = $3,150,000 - $6900 = $3,143,100

b) New Bond

Interest on new bonds = $60,000,000 \times 6% = $3,600,000

Less: Tax @30% = $3,600,000 \times 0.3 = $1,080,000

After tax interest = $3600,000 - $1,080,000 = $2,520,000

flotation cost to be amortized per annum on new bonds = $360,000 \div 12 = $30,000

Tax saving on flotation cost of new bonds = $30,000 \times 0.3 = ($9,000)

Annual after tax cost payment under new bond (b) = $2,511,000

Annual cashflows savings (a) - (b) = $632,100

Analysis of proposed refunding

NPV of the proposal

Present value of annual cashflow savings = $632,100 \times PvAF(4.2%,12y) = 9.2771 \times $632100 = $5,864,054.91

Less: Initial investment = $3,907,800.

NPV = $1,956,254.91

Since NPV is positive.Bond refunding proposal can be recommended.